complete the ordered pairs for the equation 2x+y=10
(5, ),( ,10),(,2) (0,)

Answer by funmath(2873)

complete the ordered pairs for the equation 2x+y=10
(5, ),( ,10),(,2) (0,)
Ordered pairs are in this order (x,y)
For (5, ), let x=5 and solve for y:
2(5)+y=10
10+y=10
-10+10+y10-10
y=0 The ordered pair is (5,0)
:
For ( ,10) let y=10 and solve for x:
2x+(10)=10
2x+10-10=10-10
2x=0
2x/2=0/2
x=0 The ordered pair is (0,10)
:
For ( ,2), let y=2 and solve for x:
2x+2=10
2x+2-2=10-2
2x=8
2x/2=8/2
x=4 The ordered pair is (4,2)
:
For (0, ) let x=0 and solve for y:
2(0)+y=10
0+y=10
y=10 The ordered pair is (0,10) <--we knew that already.
Happy Calculating!!!